# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v2

@@ 5 When will Personalization Improve Communication Efficiency
As a standard argument about the convergence rate (T Dinh et al., 2020; Li et al., 2021; Chen et al., 2023b), we assume the loss function $\ell$ is $L$-smooth.
Assumption 4 (Smoothness).
Based on Assumption 4 and the new notations in (7) and (8), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in B.1.
@@
Under Assumption 2 and 4, $F_{i}$ is $\mu_{g}$-strongly convex and $L_{g}$-smooth, with $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ and $L_{g}=\frac{\lambda L}{\lambda+L}$, each $h_{i}$ is $\mu_{\ell}$-strongly convex and $L_{\ell}$-smooth, with $\mu_{\ell}=\mu+\lambda$ and $L_{\ell}=L+\lambda$, and the mapping $\bm{w}_{\star}^{(i)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L_{w}$-Lipschitz with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
@@
where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$. The overall procedure is summarized in Alg. 1.
for $t=1,\dots,T$ do
       The server sends ${\bm{w}}_{t}^{(g)}$ to client $i$, $\forall i\in[m]$
       Set ${\bm{w}}_{t,0}^{(i)}={\bm{w}}_{t-1,K}^{(i)}$;
       for $k=0,\dots,K-1$ do
Algorithm 1 One-stage FedProx
@@
Alg. 1 is a customization of the general bi-level problem (Ji et al., 2022). While the algorithm design is not new, our contribution lies in analyzing how the personalization degree influences the trade-off between communication and local computation leveraging the FedProx problem structure. For simplicity, we focus on the full gradient update in our analysis; both the algorithm and the analysis can be extended to the stochastic setting with mini-batch gradients, with partial participation and incorporating variance reduction techniques.
@@
Suppose Assumptions 2 and 4 hold. Let $\{{\bm{w}}_{t}^{(g)}\}_{t\geq 0}$ and $\{{\bm{w}}_{t,K}^{(i)}\}_{t\geq 0}$ be the sequence generated by Algorithm 1 with $\gamma<1/L_{g}$, $\eta\leq 1/L_{\ell}$, and the inner loop iteration number satisfying
@@
In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\hat{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\hat{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
@@ 6 Simulation
In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions. We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3). Detailed descriptions of the experimental setup are available in Section 6.1.
@@ 6.1 Experimental Details
Real Dataset. We use the MNIST dataset for the real dataset analysis. We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using a CNN classifier. Following Li et al. (2020a), to impose statistical heterogeneity, we distribute the MNIST data across clients in the way that each client only has access to $s\in\{2,4,6,8,10\}$ classes. The fewer classes each client has access to, the higher the statistical heterogeneity. To be consistent with our theory, for both the synthetic and real datasets, we sample 2,000 samples and distribute them evenly among 10 clients, and all clients fully participate in each communication round. Additional results with 100 and 1,000 clients under partial participation are provided in Appendix C.
Selection of $\lambda$. For the logistic regression on the synthetic dataset, we selected three values for $\lambda$: small $\lambda=0.02$, medium $\lambda=0.1$, and large $\lambda=0.5$. For logistic regression on the MNIST dataset, we used small $\lambda=0.5$, medium $\lambda=1.5$, and large $\lambda=2.5$. For CNN on the MNIST dataset, we used small $\lambda=0.2$, medium $\lambda=0.5$, and large $\lambda=2.5$. Such a choice is to make sure that each line in the figure is clearly separable. Additionally, we implemented an adaptive choice of $\lambda=\frac{\rho}{\sqrt{n}R}\mathbb{I}\{R\leq n^{-1/2}\}+\frac{\rho^{2}}{nR^{2}}\mathbb{I}\{R>n^{-1/2}\}$ following the formula outlined in Theorem 1, with $\rho$ set to 2. The value of $\rho$ was determined through grid search over the range $[1,10]$ to ensure optimal performance.
Selection of Step Size. For the synthetic dataset, aligned with Theorem 2, we set the global step size $\gamma=\frac{\lambda+L}{\lambda L}$ and the local step size $\eta=\frac{1}{L+\lambda}$. The smoothness constant $L$ is computed as the upper bound of the $L$-smoothness constant for logistic regression, specifically $4^{-1}\Lambda_{\text{max}}$, where $\Lambda_{\text{max}}$ denotes the largest eigenvalue of the Gram matrix $\bm{X}^{\top}\bm{X}$. The local step size for the synthetic dataset was further tuned through grid search in $\{0.001,0.01,0.1\}$. For the MNIST dataset, the global learning rate was set to $\frac{1}{\lambda}$, while the local learning rate was set to 0.01, chosen via grid search. For both GlobalTrain and LocalTrain training baselines, the learning rate was also chosen via grid search within the range $\{0.001,0.01,0.1\}$. LocalTrain is implemented using a local (S)GD, and GlobalTrain is implemented using FedAvg (McMahan et al., 2017).
Selection of Hyperparameters. For the MNIST dataset, we implemented logistic regression using the SGD solver with 5 epochs, a batch size of 32, and 20 total runs. For the CNN model on the MNIST dataset, the same hyperparameter setup was used, except the batch size was reduced to 16 to ensure stable convergence.
Optimizer. For the synthetic dataset, following the analysis in the paper, the loss function is optimized using gradient descent (GD). For the MNIST dataset, all models are implemented in PyTorch, and optimization is performed using the stochastic gradient descent (SGD) solver.
@@ 6.2 Statistical Aspect: Balancing Local and Global Learning
From Figure 1(a), in a controlled convex setting, we observe that FedProx consistently interpolates between GlobalTrain and LocalTrain. As $\lambda$ increases, FedProx transitions from behaving similarly to LocalTrain towards GlobalTrain. When statistical heterogeneity is low, both GlobalTrain and FedProx with larger $\lambda$ values achieve better statistical performance. However, as heterogeneity increases, LocalTrain and FedProx with smaller $\lambda$ values show better results. The figure also demonstrates that with an adaptive choice of $\lambda$ outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity. See more results and discussion in Appendix C. Note that as statistical heterogeneity increases, LocalTrain experiences a slight rise in error due to the increased difficulty of the statistical problem. However, it remains the least affected by heterogeneity since it relies solely on local data. In Figure 1(b) and 1(c), for the MNIST dataset, which includes both convex logistic regression and non-convex CNN settings, we observe similar patterns. FedProx again strikes a balance between GlobalTrain and LocalTrain, adapting its performance as $\lambda$ changes. While GlobalTrain performs best in this scenario due to the relatively weak statistical heterogeneity, FedProx, with increasing $\lambda$, moves closer to GlobalTrain performance. In Figure 4 in Appendix C, we also evaluate the local model’s statistical error and observe that as statistical heterogeneity increases, FedProx with smaller $\lambda$ values can surpass GlobalTrain.
@@ 7 Conclusion
In this paper, we analyzed the effect of personalization in FedProx. With an adaptive choice of personalization degree, FedProx achieves a nearly minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training. As a minor contribution, we provably show that personalization reduces communication costs without increasing the computation overhead. The results are validated through extensive experiments.
Appendix A Proof of statistical convergence
@@ A.1 Discussion on Assumption 1
Under Assumptions 2 and 4 as we assume, Assumption 1 about statistical heterogeneity would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020a), parameter difference (Chen et al., 2023b) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
A.2 Proof of global statistical Error Bound
In this part, we will establish the statistical error bound for the global model with fewer assumptions: firstly, in the objective (1), we assume the vector $\mathcal{P}=(p_{1},\cdots,p_{m})$ just under the constraint: $p_{i}\geq 0,\forall i\in[m]$ and $\sum_{i\in[m]}p_{i}=1$; secondly, we assume the data size of the client $i$ is $n_{i}$ for any $i\in[m]$. For convenience, we neglect the notation of taking an expectation over all data during the inducing process.
First, we will prove Lemma 1, which is a crucial part to prompt the analysis of the global statistical error bound.
Proof.
We introduce some basic results from the optimality condition of the objective (1), which will be utilized later.

### v3

@@ 5 When will Personalization Improve Communication Efficiency
Based on Assumption 2 and the new notations in (8) and (9), the following lemma describes the optimizing properties for $F_{i}$, $h_{i}$ and ${\bm{w}}_{\star}^{(i)}$. The proof can be found in Appendix B.2.
@@
Under Assumption 3 and 2, $F_{i}$ is $\mu_{g}$-strongly convex and $L_{g}$-smooth, with $\mu_{g}=\frac{\lambda\mu}{\lambda+\mu}$ and $L_{g}=\frac{\lambda L}{\lambda+L}$, each $h_{i}$ is $\mu_{\ell}$-strongly convex and $L_{\ell}$-smooth, with $\mu_{\ell}=\mu+\lambda$ and $L_{\ell}=L+\lambda$, and the mapping $\bm{w}_{\star}^{(i)}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is $L_{w}$-Lipschitz with $L_{w}=\frac{\lambda}{\lambda+\mu}$.
@@
where $\nabla h_{i}({\bm{w}}_{t,k}^{(i)},{\bm{w}}_{t}^{(g)})$, for notation simplicity, denotes the partial gradient of $h_{i}$ with respect to ${\bm{w}}^{(i)}$. The overall procedure is summarized in Alg. 1, see Appendix B.1.
@@
Alg. 1 is a customization of the general bi-level problem (Ji et al., 2022). While the algorithm design is not new, our contribution lies in analyzing how the personalization degree influences the trade-off between communication and local computation leveraging the FedProx problem structure. Extending to noisy settings to consider stochastic gradient or variance reduction would be interesting future work, which requires additional technical effort.
@@
Suppose Assumptions 3 and 2 hold. Let $\{{\bm{w}}_{t}^{(g)}\}_{t\geq 0}$ and $\{{\bm{w}}_{t,K}^{(i)}\}_{t\geq 0}$ be the sequence generated by Algorithm 1 with $\gamma<1/L_{g}$, $\eta\leq 1/L_{\ell}$, and the inner loop iteration number satisfying
@@
In the same setting as Theorem 2, if choosing the step size $\eta=1/L_{\ell}=(\lambda+L)^{-1}$, $\gamma=1/(2L_{g})=\frac{\lambda+L}{2\lambda L}$, and the inner loop iteration number $K=\tilde{\mathcal{O}}\left(\frac{\lambda+L}{\lambda+\mu}\right)$, then the communication complexity and computation complexity for Algorithm 1 to find an $\varepsilon$-solution (i.e., $\|{\bm{w}}_{T}^{(g)}-\tilde{{\bm{w}}}^{(g)}\|^{2}\leq\varepsilon$ and $\|{\bm{w}}_{T,0}^{(i)}-\tilde{{\bm{w}}}^{(i)}\|^{2}\leq\varepsilon$) are as follows ($\kappa:=L/\mu$):
@@ 6 Simulation
In this section, we provide empirical validation for our theoretical results, evaluating FedProx on both synthetic and real datasets with convex and non-convex loss functions. We first examine the effect of personalization on statistical accuracy under different levels of heterogeneity (Section 6.2), followed by an analysis of its impact on communication efficiency (Section 6.3). Detailed descriptions of the experimental setup are available in Section 6.1 and Appendix C, and the complete anonymized codebase is accessible at https://anonymous.4open.science/r/fedprox-bilevel.
@@ 6.1 Experimental Details
Real Dataset. We use the MNIST, EMNIST, and CIFAR10 datasets for the real dataset analysis. We first formulate the problem as a convex classification task using multinomial logistic regression, then test the generalization of our theory to a more general, non-convex scenario by conducting experiments using different CNN classifiers. Following Li et al. (2020a), to impose statistical heterogeneity, we distribute the data across clients in the way that each client only has access to a fixed number of classes. The fewer classes each client has access to, the higher the statistical heterogeneity. The detail of data processing to set data heterogeneity for each dataset can be found in Appendix C.
Implementation and Evaluation. For each setup, we evaluate the FedProx algorithm with three different choices of $\lambda$: small, medium, and large, with specific values varying based on the dataset (details provided in Appendix C). For the synthetic dataset, in alignment with theorem 2, we implement Algorithm 1 using a global step size $\gamma=(\lambda+L)/(\lambda L)$ and a local step size $\eta=(L+\lambda)^{-1}$. For the real dataset, since $L$ is unknown, we implement the same algorithm with a global step size $\gamma$ set to $1/(\lambda)$, while the local step size is determined via grid search. In terms of evaluation, for the synthetic dataset, we evaluate by tracking the ground-truth models and measuring error as the distance from these true models. For the real dataset, we report training loss and testing accuracy. Additional details, including hyperparameter setting and evaluation metric definition, can be found in Appendix C.
@@ 6.2 Statistical Aspect: Balancing Local and Global Learning
From Figure 1(a), in a controlled convex setting, we observe that FedProx consistently interpolates between GlobalTrain and LocalTrain. As $\lambda$ increases, FedProx transitions from behaving similarly to LocalTrain towards GlobalTrain. When statistical heterogeneity is low, both GlobalTrain and FedProx with larger $\lambda$ values achieve better statistical performance. However, as heterogeneity increases, LocalTrain and FedProx with smaller $\lambda$ values show better results. The figure also demonstrates that with an adaptive choice of $\lambda$ in theoretical analysis outlined in Theorem 1, FedProx consistently achieves a low error across varying levels of heterogeneity. See more results and discussion in Appendix C. Note that as statistical heterogeneity increases, LocalTrain experiences a slight rise in error due to the increased difficulty of the statistical problem. However, it remains the least affected by heterogeneity since it relies solely on local data. In Figure 1(b) and 1(c), for the MNIST dataset, which includes both convex logistic regression and non-convex CNN settings, we observe similar patterns. FedProx again strikes a balance between GlobalTrain and LocalTrain, adapting its performance as $\lambda$ changes. While GlobalTrain performs best in this scenario due to the relatively weak statistical heterogeneity, FedProx, with increasing $\lambda$, moves closer to GlobalTrain performance. In Figure 4 in Appendix C, we also evaluate the local model’s statistical error and observe that as statistical heterogeneity increases, FedProx with smaller $\lambda$ values can surpass GlobalTrain.
@@ 7 Conclusion
In this paper, we analyzed of the effect of personalization in FedProx. With a theoretically-informed adaptive choice of personalization degree, FedProx achieves a minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training. As a minor contribution, we provably show that personalization reduces communication costs without increasing the computation overhead. The results are validated through extensive experiments.
Appendix A Proof of statistical converngence
@@
And we also denote
$\bm{\delta}_{\star}^{(i)}:={{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}},\quad\tilde{\bm{\delta}}^{(i)}:={\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}},$ (17)
where $\bm{\delta}_{\star}^{(i)}$ measures the difference between the true global model ${{\bm{w}}_{\star}^{(g)}}$ and the true local model ${{\bm{w}}_{\star}^{(i)}}$, and $\tilde{\bm{\delta}}^{(i)}$ is the estimator of such a difference.
@@ A.1 Discussion on Assumption 1
Under Assumptions 2 and 3 as we assume, Assumption 1 about statistical heterogeneity would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020a), parameter difference (Chen et al., 2023b) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
A.2 Lemmas and the Proof
To facilitate the proof of the theorem, we first present two additional lemmas and provide their proof.
Lemma 4.
Under Assumption 3, for the optimal solution ${\tilde{\bm{w}}^{(g)}}$ and ${\tilde{\bm{w}}^{(i)}}$’s in Problem (1), we have
$\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|_{2}\leq\frac{2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|_{2}}{\mu+\lambda}.$
By the $\mu$-strongly convexity of $\nabla L_{i}$, we have that for any ${\bm{w}}\in\mathbb{R}^{d}$,
$L_{i}({\bm{w}},S_{i})+\frac{\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$ $\geq L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\left\langle\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i}),{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\right\rangle+\frac{\mu+\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$
$\geq L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|\left(\frac{\mu+\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|-\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|\right),$ (19)
where in the last step we used the Cauchy inequality.
In addition, notice that ${\tilde{\bm{w}}^{(i)}}$ is the optimal solution for the local model ${{\bm{w}}^{(i)}}$ in the FedProx problem (1), thus we have
${\tilde{\bm{w}}^{(i)}}=\underset{{\bm{w}}}{\operatorname{argmin}}\left\{L_{i}({\bm{w}},S_{i})+\frac{\lambda}{2}\|{\bm{w}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\right\}.$
This result together with (A.2) implies that
$L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})$ $=L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}\geq L_{i}({\tilde{\bm{w}}^{(i)}},S_{i})+\frac{\lambda}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|^{2}$
$\geq L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})+\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|\left(\frac{\mu+\lambda}{2}\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|-\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|\right).$
Therefore, since $\mu+\lambda>0$ and $\|{\tilde{\bm{w}}^{(i)}}-{\tilde{\bm{w}}^{(g)}}\|\geq 0$, rearranging the terms leads to
$\|{\tilde{\bm{w}}^{(g)}}-{\tilde{\bm{w}}^{(i)}}\|\leq\frac{2\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|}{\mu+\lambda},$ (20)
as claimed.
Lemma 5.
Under Assumption 1, 2, 3, 4, for $\tilde{\bm{w}}_{\text{GT}}$, the solution of the GlobalTrain problem in (2), we have
$\mathbb{E}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|\leq\frac{\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}+LR}{\mu/2}.$
By the optimality condition of the GlobalTrain problem in (2) and the strong convexity of $L_{i}$’s, we have
00 0 $\geq\sum_{i\in[m]}p_{i}\left(L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})-L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})\right)$
$\geq\sum_{i\in[m]}p_{i}\left\langle\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i}),\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\right\rangle+\frac{\mu}{2}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}$
$\geq-\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})\Big{\|}\Big{\|}\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\Big{\|}+\frac{\mu}{2}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}.$
Therefore, we have
$\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|$ $\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})\Big{\|}}{\mu/2}$
$\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}+\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}}{\mu/2}$ (21)
By the $L$-smoothness property of $L_{i}$’s in Assumption 2 and the bounded gradient property listed in Assumption 4, we have
$\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}\leq\sum_{i\in[m]}p_{i}L\|{{\bm{w}}_{\star}^{(g)}}-{{\bm{w}}_{\star}^{(i)}}\|\leq LR,$
$\mathbb{E}\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}=\mathbb{E}\Big{\|}\sum_{i\in[m]}\sum_{j\in[n_{i}]}\frac{p_{i}}{n_{i}}\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij})\Big{\|}\leq\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}.$
These results together with (A.2) implies
$\mathbb{E}\|\tilde{\bm{w}}_{\text{GT}}-{{\bm{w}}_{\star}^{(g)}}\|\leq\frac{\rho\sqrt{\sum_{i\in[m]}\frac{p_{i}^{2}}{n_{i}}}+LR}{\mu/2}$
as claimed.
A.3 Proof of Lemma 1 and Theorem 1
A.3.1 Proof of Lemma 1
Lemma 1 is a direct consequence of the proof of Theorem 1. To see this, from (A.3.2) and the fact that $(\mu+\lambda)^{2}\geq\mu(\mu+\lambda)$, we have
$\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}$
$\leq\frac{4(1+L/\mu)}{\mu(\mu+\lambda)}\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2},$
thus (4) holds. In addition, with the fact that $\|\tilde{\bm{\delta}}^{(i)}\|_{2}\geq 0$, from (28) we find
00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}$
$\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$
which implies the result in (5).
A.3.2 Proof of Global Statistical Error Bound in Theorem 1
Case 1: We first consider the case when $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.